Unit 1 Review:

8) Which graph of inverse of f(x)?


1
Expert's answer
2022-02-02T11:11:14-0500

For example, let "f(x)=x+3." Then "f^{-1}(x)=x-3" . The graph is as below:



The graph of "f(x)" is green while, the inverse is red